OpenStudy (anonymous):

What is the sum of the geometric sequence 1, -4, 16, ... if there are 7 terms? Choose one answer. a. -5,461 b. 5,461 c. -3,277 d. 3,277

OpenStudy (callisto):

Sum of GS \(\large = \frac{a(1-r^n)}{1-r}\) Now, a=1 , r=-4/1 = -4 , n=7 Sum of GS \(\large = \frac{(1)(1-(-4)^7)}{1-(-4)} = ...?\)